OpenStudy (carolina4567):

To open these doors, you must match the number and type of solutions for the following two functions in standard form. f(x) = x2 + 6x – 16 g(x) = x2 +6x + 1 Match the following descriptions of the solutions to each of the functions above. Hint: they each have their own match. Two real irrationals solutions Two real rationals solutions

OpenStudy (carolina4567):

Help me @mathmale @iGreen

OpenStudy (carolina4567):

SO I was thinking that F(x) has two real rationals solutions and G(x) has two real irrational solutions

OpenStudy (mathmale):

Strictly speaking, f(x) = x2 + 6x - 16 and g(x) = x2 +6x + 1 don't have solutions; it's not until you set each one equal to zero that the possibility of finding solutions arises. So: Find the solutions of f(x) = x2 + 6x - 16=0 and of g(x) = x2 +6x + 1=0.

OpenStudy (mathmale):

By the way, for clarity, please type "the square of x" as x^2, or as \[x^2\or write it by hand:

OpenStudy (mathmale):

\[x^2\]

OpenStudy (carolina4567):

Im not sure how to solve @mathmale

OpenStudy (carolina4567):

Im having a really hard time with this assigment and its the last one i need to do to finish :(

OpenStudy (mathmale):

Just what is it about the assignment that is hard for you? Put that into words, please.

OpenStudy (carolina4567):

The solving quadricts and all of that stuff: I would give it a try though: Is it like this f(x) = x2 + 6x – 16 x1= 2 x2=-3 for g(x) = x2 +6x + 1 x1= −3+22√ x2= −3−22√

OpenStudy (carolina4567):

Its -3-2 square root 2

OpenStudy (mathmale):

Do you mean the following?